OpenStudy (anonymous):

solve: -2^x = (-1/4) would the answer be x = -2 ?

OpenStudy (lgbasallote):

let's check \[-2^x = -\frac 14\] \[2^x = \frac 14\] \[2^x = 4^{-1}\] \[2^x = (2^2)^{-1}\] \[2^x = 2^{-2}\] x = -2 not sure if you're familiar with the method i just used

OpenStudy (lgbasallote):

so bottomline you're right lol

OpenStudy (anonymous):

okay, thank you :)